(b) How many grams of carbon dioxide can be produced by burning 7.16 moles of propane? Assume that oxygen is the excess reactant in this reaction.

1
Expert's answer
2022-02-17T19:19:03-0500

2 C3H8 + 9 O2 → 4 CO2 + 2 CO + 8 H2O

mole ratio= 2:4= 1:2

Moles of CO2 produced= 7.16× 2= 14.32moles

Mass= 14.32× 44= 630g